OpenStudy (blondisaurus):

Find all zeros of the function f(x)= x^2(x+3) Explain the step by step process to answer these types of questions please :(

OpenStudy (anonymous):

put x^2 (x+3)=0 find x

OpenStudy (blondisaurus):

Would you distribute the x^2 first?

OpenStudy (blondisaurus):

HELP

OpenStudy (amoodarya):

x^2=0 or x+3=0 so x^2=0 --> x=0 or x=-3

OpenStudy (blondisaurus):

Ohhh, the book mentions 0 twice. But ok i get it